OpenStudy (anonymous):

What's x^2 + 3^2=(3√3)^2 . I'm using the Pythagorean theorem is this possible?

OpenStudy (anonymous):

Yes, it is possible.

OpenStudy (anonymous):

Are you solving for \(x\) then?

OpenStudy (anonymous):

Yes, it is possible....

OpenStudy (usukidoll):

as long as the equation is in the form a^2+b^2 =c^2 you can

OpenStudy (whpalmer4):

You need to figure out what \((3\sqrt{3})^2 =\) Hint:\((3\sqrt{3})^2 = (3\sqrt{3})(3\sqrt{3}) = 3*3*\sqrt{3}*\sqrt{3} = 3*3*\sqrt{3*3}\)